if the two angles of a quadrilateral are supplementary and the third angle is 30degrees ,then find the fourth angle

 

Search Results

Videos

Grade:6

1 Answers

Arun
4 years ago
 
if two angles(A and B) are supplementary then
A + B = 180
 
now A + B + C + D = 360
 
D+ 180 + 30 = 360
 
hence D = 360 – 210 = 150
 
Regards
Arun (askIITians forum expert)